OpenStudy (anonymous):

Find the exact area of the circle with the given diameter. d = 12.6 cm A. 39.69 cm2 B.158.76 cm2 C.158.76π cm2 D. 39.69π cm2 plz help will medal

OpenStudy (ikram002p):

the area of a circle A with diameter d is denoted by \(A=(\dfrac{d}{2} )^2 \pi\) s.t \(\pi=3.14\)

OpenStudy (anonymous):

so is it d?

OpenStudy (anonymous):

Which is the closest approximation to the area of the circle? Use 3.14 or to approximate π. A. 57 mm² B. 113 mm² C. 324 mm² D. 1017 mm² plz help will give medal

OpenStudy (ikram002p):

first Qn idk which choise is it wolfram gives this http://www.wolframalpha.com/input/?i=%2812.6%2F2%29^2+*+3.14

OpenStudy (anonymous):

thanx

OpenStudy (ikram002p):

np :)

OpenStudy (anonymous):

D was right:)

OpenStudy (anonymous):

What is the exact area of a circular courtyard with a radius of 22 ft? A. 1936π ft² B. 484π ft² C. 44π ft² D. 22π ft²

OpenStudy (anonymous):

will medal